Factor the expression 4x^2- 20xy +25y^2 into a product of binomials. Please show your steps.
When I worked it out ....
I got a= 2 am I right?
Thank you in advance...

Found 2 solutions by stanbon, dfrazzetto:
Answer by stanbon(75887)

Factor the expression 4x^2- 20xy +25y^2 into a product of binomials.
--------------
= (2x-5y)(2x-5y)
===================
Cheers,
Stan H.
===================

Answer by dfrazzetto(283)

4x^2- 20xy +25y^2
this is a perfect square trinomial
4x^2 = (2x)^2; 25y^2 = (5y)^2
Factoring:
4x^2- 20xy +25y^2
(2x - 5y)(2x - 5y) = (2x - 5y)^2
Check, expand the binomial squared (using FOIL):
(2x - 5y)(2x - 5y) = 4x^2 - 10xy - 10xy + 25y^2 = 4x^2- 20xy +25y^2